OpenStudy (anonymous):

College Algebra: 3X3 Linear Systems! Can somebody work this out step-by-step with me? The problem is: "Stocks, Bonds, and a Mutual Fund: Marita invested a total of $25,000 in stocks, bonds, and a mutual fund. In one year she earned 8% on her stock investment, 10% on her bond investment, and 6% on her mutual fund, with a total return of $1,860. Unfortunately, the amount invested in the mutual fund was twice as large as the amount she invested in the bonds. How much did she invest in each?"

OpenStudy (dan815):

let x be money in stocks let y be money in bonds let z be money in mutual

OpenStudy (anonymous):

I had the second equation right, but the first and third wrong. I set it that way.

OpenStudy (anonymous):

where did you get z=2y?

OpenStudy (dan815):

x+y+z=25000 0,08*x+0.10*y+0.06*z=1860 z=2*(y)

OpenStudy (dan815):

he amount invested in the mutual fund was twice as large as the amount she invested in the bonds.

OpenStudy (anonymous):

Ohhh, okay

OpenStudy (anonymous):

So I plug 2y back into .08x+.10y+.6z=1860?

OpenStudy (dan815):

0.06*z

OpenStudy (anonymous):

Yeah that's what I meant :D

OpenStudy (dan815):

plug z=2y into both equations x+y+z=25000 0,08*x+0.10*y+0.06*z=1860

OpenStudy (dan815):

you will have an 2 by 2 linear system then, you can solve that same was as yesterday then

OpenStudy (dan815):

you will have an 2 by 2 linear system then, you can solve that same wa* as yesterday then

OpenStudy (dan815):

i give up

OpenStudy (dan815):

spelling is too hard for me obviously

OpenStudy (anonymous):

lolx it's okay Dan. So, my two linear systems are x+3y=25,000 and 0.08x+0.22y=1,860

OpenStudy (anonymous):

haha, don't worry. Everyone gets better at it eventually

OpenStudy (dan815):

yes

OpenStudy (anonymous):

Wouldn't I make x+3y=25,000 into 3y=25,000-x

OpenStudy (anonymous):

Oh wait

OpenStudy (anonymous):

wrong

OpenStudy (anonymous):

I would make x+3y=25,000 into x=25,000-3y

OpenStudy (dan815):

both work

Directrix (directrix):

Let S represent the initial investment in Stocks. Let B represent the initial investment in Bonds. Let F represent the initial investment in Mutual Funds. S + B + F = 25000 .08S + .10B + .06F = 1860 F = 2B Those are your three equations in three unknowns. I don't know if you are supposed to solve those manually or if you are allowed to use technology. Using technology, you can see what the results are at: http://www.wolframalpha.com/input/?i=.08x+%2B+.10y+%2B+.06z+%3D+1860%2C+x+%2B+y+%2B+z+%3D+25000%2C+z+%3D+2y where x is Stocks, y is Bonds, and z if Mutual Funds @Natasha.g.2013 This problem can also be done with one equation but three equations were specified in the problem statement. Answered on your first post of the problem:

OpenStudy (anonymous):

Well I plugged in x=25,000-3y into 0.08x+0.22y=1,860, solved it, and got y=7000

Directrix (directrix):

I agree. @Natasha.g.2013

OpenStudy (anonymous):

Well, thank you for that Directrix, but I have to write everything out because I won't have a computer do my final exam.

OpenStudy (anonymous):

I have the answers in the back of the textbook, but I want to know how to solve word problems step by step.

OpenStudy (anonymous):

Thank you to everyone who helped. :)

OpenStudy (dan815):

np!